Mass of crucible+lid+Mg= 10.90



Crucible+lid= 10.03



Mass of MgO formed - Mass of crucible= 10.90-10.03= 0.87



Calculate:



Mass of MgO



Amount of Mg in the compound



Mass of O2 in the compound



Amount of oxygen in mole



Simple number ratio to oxygen to Mg



Empirical formula of the compound.




1
Expert's answer
2022-01-25T00:50:01-0500

All information not given.


Suppose mass of crucible + lid + MgO"=11.47"

Mass of Mg "=10.90-10.03"

"=0.87"


Mass of MgO"=11.47-10.03"

"=1.44"


Mass of O"=1.44-0.87=0.57"


Number Of moles of Mg"=\frac{0.87}{24.305}=0.035795"


Number of moles of O"=\frac{0.57}{16}=0.035625"


Oxygen to Magnesium "=1:1"


Empirical formula "=" MgO